OpenStudy (anonymous):

how do i graph (y=1/3x-2

OpenStudy (anonymous):

put a dot at y = -2. From that dot go up one unit and then three units to the right and put another dot. From that dot go up one unit and then go three units to the right and put a dot. Connect the dots.

OpenStudy (saifoo.khan):

slope = 1/3 y intercept = -2/

OpenStudy (electrochika):

start at -2 go 1 over 3 up

OpenStudy (electrochika):



OpenStudy (anonymous):

y-int is (0,-2) then pick values of x that are mutliples of 3 since slope is m=1/3 (easy arithmetic) see attached; I chose x=3, 6 as examples